OpenStudy (anonymous):

Find the slope of the line that contains the following pair of points. (2, 4), (0, 8) A. -2 B. -1/2 C. 1/6 D. 6

OpenStudy (uri):

m=(y2-y1)/x2-x1)

OpenStudy (uri):

m=8-4/0-2 m=4/-2 m=-2

OpenStudy (khaakenajaf):

-2 answer is a